# Can Test-time Computation Mitigate Memorization Bias in Neural Symbolic Regression?

### v1

@@ Abstract
Symbolic regression aims to discover mathematical equations that fit given numerical data. It has been applied in various fields of scientific research, such as producing human-readable expressions that explain physical phenomena. Recently, Neural symbolic regression (NSR) methods that involve Transformers pre-trained on large-scale synthetic datasets have gained attention. While these methods offer advantages such as short inference time, they suffer from low performance, particularly when the number of input variables is large. In this study, we hypothesized that this limitation stems from the memorization bias of Transformers in symbolic regression. We conducted a quantitative evaluation of this bias in Transformers using a synthetic dataset and found that Transformers rarely generate expressions not present in the training data. Additional theoretical analysis reveals that this bias arises from the Transformer’s inability to construct expressions compositionally while verifying their numerical validity. We finally examined if tailoring test-time strategies can lead to reduced memorization bias and better performance. We empirically demonstrate that providing additional information to the model at test time can significantly mitigate memorization bias. On the other hand, we also find that reducing memorization bias does not necessarily correlate with improved performance. These findings contribute to a deeper understanding of the limitations of NSR approaches and offer a foundation for designing more robust, generalizable symbolic regression methods. Code is available at https://github.com/Shun-0922/Mem-Bias-NSR.
@@ 1 Introduction
Discovering underlying equations from collected experimental data is a crucial process in many fields of scientific research. Symbolic regression is a branch of regression analysis that seeks to automatically identify underlying mathematical expressions. In contrast to methods that model data without explicit mathematical expressions, symbolic regression offers advantages in terms of interpretability and generalizability. This is because the outputs of symbolic regression are usually compact, human-readable equations, making them less susceptible to overfitting. However, symbolic regression is a challenging task due to its vast search space; the number of possible mathematical expressions grows exponentially with expression length or the number of input variables. Applications for symbolic regression span various fields of scientific research such as physics [27], materials science [35], and weather forecasting [1].
@@
Various methods for symbolic regression have been proposed in recent years. Traditionally, approaches based on genetic programming (GP) [11] have been employed to solve symbolic regression. These methods tend to be computationally expensive because they generate each expression entirely from scratch. To mitigate this inefficiency, a research direction called neural symbolic regression (NSR) has emerged. NSR methods leverage encoder-decoder Transformer architectures [32] pre-trained on large-scale synthetic datasets [4, 30]. Since a single forward pass through the Transformer suffices to output a candidate equation, NSR models can generate solutions far more quickly than GP-based approaches. However, NSR methods often falls short in terms of numerical accuracy, with particularly poor performance when the number of input variables is large [10, 3].
@@
This study aims to uncover the underlying cause of this drawback and explore methods to alleviate it. Our hypothesis is that, in NSR methods, naively using a Transformer for inference leads to memorization bias, meaning that models struggle to generate novel expressions not seen during training. Given that the expressions in the training data typically represent only a small subset of the full space of possible expressions, our hypothesis implies that standard NSR methods operate within a significantly constrained search space. We investigated this hypothesis in NeSymReS [4], a pioneering work in NSR models. We created a synthetic training dataset with a restricted set of operators to analyze NeSymReS’s memorization bias, and found that it fails to produce expressions outside the training data. Prior work has highlighted NSR methods’ limited ability to extrapolate to unseen numerical domains—e.g., models trained on data whose input variable $x$ lies in the interval $[-1,1]$ often fail when evaluated on inputs from the wider interval $[-2,2]$ [16, 23]. However, the limitation we identify is even more fundamental: NSR models often fail to generalize even within their training domain. This work is the first to show that standard NSR models primarily memorize training expressions instead of composing familiar components into genuinely novel formulas.
To understand why NeSymReS cannot generate unseen expressions, we conducted further theoretical analysis. We formulate a simplified symbolic regression task to evaluate whether Transformers possess the ability to compositionally generate mathematical expressions. We showed that the Transformer fails to solve such tasks due to its inability to numerically verify the expressions it generates. Our analysis demonstrates that NSR models’ memorization bias is inherently rooted in their architectural structure.
@@
Towards the end of this paper, we explore methodologies to mitigate the memorization bias of standard NSR models and improve numerical accuracy. We focus particularly on test-time strategies and investigate how they affect memorization bias and numerical accuracy. We compared three strategies: decoding with a large beam size, decoding using MCTS, and providing verification feedback at the subtree level. The last strategy is a new method that we propose, which we refer to as neural symbolic regression guided by verified subtrees (NSR-gvs). We find that providing new information to the model during test-time leads to generating expressions beyond the training dataset. However, we also identified cases where memorization bias was mitigated but numerical accuracy decreased, as well as cases where numerical performance improved despite little reduction in memorization bias. We conclude this paper by discussing the underlying causes of these differences across varying types of test-time strategies.
@@
  • We quantitatively demonstrate, under controlled settings, that naively applying a Transformer to symbolic regression leads to memorization bias.
  • To account for the observed memorization bias, we conducted a theoretical analysis and formally demonstrate that Transformers lack the ability to compositionally generate expressions while accounting for numerical data.
  • We compare varying test-time computing strategies and analyze how such strategies affect memorization bias and numerical accuracy.
@@ 2 Related Work
Several approaches to symbolic regression exist, such as GP, brute force algorithms, reinforcement learning, and NSR. Since our study focuses on analyzing and improving NSR methods, we describe NSR in detail in this section, and provide explanation for other symbolic regression methods in Section F.
Traditional symbolic regression methods such as GP generate each equation from scratch, resulting in long inference times, with equation generation potentially taking hours. In order to achieve a shorter inference time, studies such as NeSymReS [4] and SymbolicGPT [30] carried out large scale pre-training of Transformers. In these studies, an artificial dataset consisting of millions of randomly generated equations was used for training. These methods, often categorized as NSR, can generate an expression in just a few seconds, significantly reducing inference time compared with other approaches.
In the recent years, a number of studies, summarized in Table 1, have focused on enhancing NSR methods. Studies such as [10] and [31] proposed an end-to-end approach using a Transformer model to directly predict full mathematical expressions including constants, whereas previous methods followed a two-step procedure where constant fitting had to be done separately. Lalande et al. [12] analyzed several different architectures to find the suitable encoder architecture for NSR. Shojaee et al. [23] focused on improving the decoding strategy for NSR, incorporating the MCTS algorithm during the generation of expressions. In their study, Li et al. [16] trained a Transformer model to imitate the process of improving mathematical formulas, as performed in the reinforcement learning-based approach proposed by Mundhenk et al. [20]. Bendinelli et al. [3] proposed a model called NSRwH that enables incorporating prior knowledge, which is often available during application in scientific research. For example, scientists may anticipate symmetries between variables or expect certain partial expressions to appear in the mathematical laws governing the data. NSRwH adds a dedicated encoder that processes such prior knowledge, allowing the model to generate expressions that are consistent with both the numerical data and the provided prior knowledge.
@@
NSR Methods Automatic Direct Information Assessing the
Training Data Constant Added During Memorization Bias
Generation Prediction Test-time
Biggio et al. [4] - - -
Kamienny et al. [10] - -
Shojaee et al. [23] MCTS Feedback -
Li et al. [16] - - Historical Context -
Bendinelli et al. [3] - Prior Knowledge -
@@ 3 Problem Formulation
We formalize NSR as the problem of learning a parameterized symbolic regressor $S_{\boldsymbol{\theta}}$ that maps a numerical dataset $\mathcal{D}$ to a symbolic expression $\hat{e}=S_{\boldsymbol{\theta}}(\mathcal{D})$. The learning algorithm is formulated as minimizing a loss that measures how well $\hat{e}$ matches the ground-truth expression $e^{\ast}$ underlying $\mathcal{D}$. In this section, we specify how synthetic training pairs $(e^{\ast},\mathcal{D})$ are generated in NeSymReS [4], since it is the foundational work underlying our research.
@@ 3.1 Synthetic Expression Distribution
[…]
Let $\mathcal{E}$ be an expression space and $p_{\mathcal{E}}$ be the generator of symbolic expressions employed in NeSymReS [4]. We also denote by $p_{\mathrm{Tree}}$ the generator of unary-binary trees introduced by Lample and Charton [13]. We write $e^{\ast}\sim p_{\mathcal{E}}$ for the following procedure.
@@ 3.2 Synthetic Dataset Generation
$\mathcal{D}\;=\;\{\,(\mathbf{x}_{i},\,y_{i})\,\}_{i=1}^{n},\qquad x_{ij}\;\sim\;\mathcal{U}([x_{\min,j},x_{\max,j}])\quad\text{for}\quad j=1,2,\ldots,d,\qquad y_{i}\;=\;e^{\ast}(\mathbf{x}_{i}).$
@@
Note that, in practice, the training dataset is the collection of $e_{\mathrm{templ}}$, and both $e^{\ast}$ and $\mathcal{D}$ are generated dynamically during training. Further details concerning the work of NeSymReS [4] are described in Appendix A.
4 Exploring Memorization Bias in NSR
In this section, we first explore NeSymReS’s memorization bias in a controlled setting. We further present theoretical analysis to explain the empirical results.
4.1 Method
We first constructed a simplified training dataset consisting of $100$K equations. The dataset size was kept relatively small due to the computational cost required to verify whether each generated expression already appears in the training set. The allowed operators were add, sub, sin, cos, tan, and exp, with up to $5$ independent variables per equation. Since this dataset is smaller than that used in the original NeSymReS setting, we also restricted the operator set to further reduce the search space. We then trained a NeSymReS model on this dataset for $1,000$ epochs.
As outlined in Section 3, the training dataset for NeSymReS comprises multiple instances of $e_{\mathrm{templ}}$, each of which is an expression tree without numerical constants. Accordingly, we assess the novelty of output expressions at the level of tree structure while ignoring numerical constants.

### v2

# Can Test-time Computation Mitigate Reproduction Bias in Neural Symbolic Regression?

@@ Abstract
Mathematical expressions play a central role in scientific discovery. Symbolic regression aims to automatically discover such expressions from given numerical data. Recently, Neural symbolic regression (NSR) methods that involve Transformers pre-trained on synthetic datasets have gained attention for their fast inference, but they often perform poorly, especially with many input variables. In this study, we analyze NSR from both theoretical and empirical perspectives and show that (1) naive token-by-token generation is ill-suited for NSR, as Transformers cannot compositionally generate tokens while validating numerical consistency, and (2) the search space of NSR methods is greatly restricted due to reproduction bias, where the majority of generated expressions are merely copied from the training data. We further examine whether tailored test-time strategies can reduce reproduction bias and show that providing additional information at test time effectively mitigates it. These findings contribute to a deeper understanding of the limitation of NSR approaches and provide guidance for designing more robust and generalizable methods. Code is available at https://github.com/Shun-0922/Mem-Bias-NSR.
Machine Learning, ICML
@@ 1 Introduction
Discovering underlying equations from collected experimental data is a crucial process in many fields of scientific research. Symbolic regression is a branch of regression analysis that seeks to automatically identify the underlying mathematical expressions. In contrast to methods that model data without explicit mathematical expressions, symbolic regression offers advantages in terms of interpretability and generalizability. This is because the outputs of symbolic regression are usually compact, human-readable equations, making them less susceptible to overfitting. However, symbolic regression is a challenging task due to its vast search space; the number of possible mathematical expressions grows exponentially with expression length or the number of input variables. Applications for symbolic regression span various fields of scientific research such as physics (Tenachi et al., 2023), materials science (Wang et al., 2019), and weather forecasting (Abdellaoui and Mehrkanoon, 2021).
@@
Various methods for symbolic regression have been proposed in recent years. Traditionally, approaches based on genetic programming (GP) (Koza, 1994) have been used to solve symbolic regression. These methods tend to be computationally expensive because they generate each expression entirely from scratch. To mitigate this inefficiency, a research direction called neural symbolic regression (NSR) has emerged. NSR methods leverage encoder-decoder Transformer architectures (Vaswani et al., 2017) pre-trained on large-scale synthetic datasets (Biggio et al., 2021; Valipour et al., 2021). NSR methods generate expressions similar to natural language processing tasks, where expressions are generated token-by-token in an auto-regressive manner. Since a single forward pass through the Transformer suffices to output a mathematical token (e.g., $x_{1},\sin,+$), NSR models can generate solutions far more quickly than GP-based approaches. However, NSR methods often fall short in terms of numerical accuracy, with particularly poor performance when the number of input variables is large (Kamienny et al., 2022; Bendinelli et al., 2023). This study aims to uncover the underlying cause of this drawback and explore methods to alleviate it.
Our analysis begins by examining the mechanisms a Transformer relies on to select the next token during expression generation. Specifically, we theoretically analyze the limitations Transformers face when generating mathematical expressions. A natural and ideal token-generation strategy would be to select, in a compositional manner at each step, the token that is most appropriate in light of the numerical data. However, by using circuit complexity theory, we show that Transformers fail to generate expressions in such ways; they cannot compositionally generate mathematical expressions while taking numerical data into account. In other words, while Transformers generate tokens sequentially, they are not able to carry out meaningful processing at every step of token generation, suggesting that naive token-by-token generation is not particularly suitable for NSR. To illustrate this limitation, consider a situation where a Transformer has generated an expression up to $x_{1}^{2}+\sin(x_{2})+$. Our analysis implies that Transformers are unable to internally compute which leaf token (e.g., $x_{1},x_{2},x_{3},\ldots$) would be the most appropriate given the input numerical data. The result indicates that in practice, Transformers generate expressions by some alternative, more coarse-grained mechanism instead of generating tokens in a compositional manner.
@@
We next investigated how NSR methods actually generate expressions under empirical conditions. We hypothesize that, in NSR methods, naively using a Transformer for inference leads to reproduction bias, meaning that models struggle to generate novel expressions not seen during training and instead tend to generate expressions copied from the training data. Given that the expressions in the training data typically represent only a small subset of the full space of possible expressions, our hypothesis implies that standard NSR methods operate within a significantly constrained search space. We investigated this hypothesis in NSR methods such as NeSymReS (Biggio et al., 2021), a pioneering work in NSR models. We found that the majority of expressions generated by Transformers are expressions that were included in the training dataset, which supports our hypothesis of reproduction bias. Prior work has highlighted NSR methods’ limited generalizability with respect to the range of numerical data—e.g., models trained on data whose input variable $x$ lies in the interval [1,1][-1,1] often fail when evaluated on inputs from the wider interval [2,2][-2,2] (Li et al., 2024; Shojaee et al., 2023). However, the reproduction bias that we identify is orthogonal to this phenomenon and represents an even more fundamental limitation: NSR models often fail to generalize even within their training domain. This work is the first to show that standard NSR models primarily copy training expressions instead of composing familiar components into genuinely novel formulas.
@@
Towards the end of this paper, we explore methodologies to mitigate the reproduction bias of standard NSR models and improve numerical accuracy. Since the theoretical analysis suggested that naive auto-regressive generation may be inappropriate for NSR, we focus particularly on test-time strategies and investigate how they affect reproduction bias and numerical accuracy. We compare three strategies: decoding with a large beam size, decoding using MCTS, and providing verification feedback at the subtree level. The last strategy is a new method that we propose, which we refer to as neural symbolic regression guided by verified subtrees (NSR-gvs). We found that providing new information to the model during test-time leads to generating expressions beyond the training dataset. However, we also identify cases where reproduction bias was mitigated but numerical accuracy decreased, as well as cases where numerical performance improved despite little alleviation in reproduction bias.
@@
  • We formally show that Transformers lack the ability to compositionally generate expressions while accounting for numerical data.
  • We empirically demonstrate, under various settings, that naively applying a Transformer to symbolic regression leads to reproduction bias.
  • We compare varying test-time computing strategies and analyze how such strategies affect reproduction bias and numerical accuracy.
@@ 2 Related Work
NSR Methods Automatic Training Direct Constant Information Added Assessing
Data Generation Prediction During Test-time Reproduction Bias
Biggio et al. (2021) - - -
Kamienny et al. (2022) - -
Shojaee et al. (2023) MCTS Feedback -
Li et al. (2024) - - Historical Context -
Bendinelli et al. (2023) - Prior Knowledge -
@@
Several approaches to symbolic regression exist, such as GP, brute-force algorithms, reinforcement learning, and NSR. Since our study focuses on analyzing and improving NSR methods, we mainly describe NSR in detail in this section, and explain other symbolic regression methods in Appendix E.
Traditional symbolic regression methods such as GP generate each equation from scratch, resulting in long inference times, with equation generation potentially taking hours. In order to achieve a shorter inference time, studies such as NeSymReS (Biggio et al., 2021) and SymbolicGPT (Valipour et al., 2021) carried out large scale pre-training of Transformers. In these studies, an artificial dataset consisting of millions of randomly generated equations was used for training. These methods, often categorized as NSR, can generate an expression in just a few seconds, significantly reducing inference time compared with other approaches.
In recent years, several studies, summarized in Table 1, have focused on improving NSR methods. Studies such as (Kamienny et al., 2022) and (Vastl et al., 2024) proposed an end-to-end approach using a Transformer model to directly predict full mathematical expressions including constants, whereas previous methods followed a two-step procedure where constant fitting had to be done separately. Lalande et al. (2023) analyzed several different architectures to find a suitable encoder architecture for NSR. Shojaee et al. (2023) focused on improving the decoding strategy for NSR, incorporating the MCTS algorithm during the generation of expressions. In their study, Li et al. (2024) trained a Transformer model to imitate the process of improving mathematical formulas, as performed in the reinforcement learning-based approach proposed by Mundhenk et al. (2021). Bendinelli et al. (2023) proposed a model called NSRwH that enables incorporating prior knowledge, which is often available during application in scientific research. For example, scientists may anticipate symmetries between variables or expect certain partial expressions to appear in the mathematical laws governing the data. NSRwH adds a dedicated encoder that processes such prior knowledge, allowing the model to generate expressions that are consistent with both the numerical data and the prior knowledge provided.
More recently, there has been growing research on methods that iteratively refine mathematical expressions, most of which rely on large language models (LLMs) rather than pre-trained Transformer models. These methods are similar to NSR-gvs, one of the test-time computation approaches considered in this paper, in that they iteratively improve their outputs by incorporating feedback from the generated expressions. In (Merler et al., 2024) and (Sharlin and Josephson, 2024), the authors introduce approaches in which a base equation structure is generated using LLMs, and the equation is subsequently improved iteratively by receiving feedback from external numerical solvers. While Shojaee et al. (2023) follow a similar methodology, it incorporates supplementary descriptions regarding the variables in the prompt, facilitating more effective use of the LLM’s scientific knowledge. Grayeli et al. (2024) proposed a method that uses an LLM to identify “concepts” representing features of high-performing expressions and leverages them to further evolve a set of equations. Zhang et al. (2025) introduces an iterative algorithm that replaces features of suboptimal expressions at each step while incorporating relevant expressions as needed. Pre-training-free methods described above may have the potential to address some of the limitations of conventional NSR approaches. On the other hand, using a pre-trained Transformer, as opposed to an LLM, offers certain advantages similar to those of small language models (SLMs), such as keeping the model size manageable and enhancing domain specificity through careful design of the training data. For these reasons, we believe that conducting a deeper analysis of pre-trained Transformer-based NSR and exploring ways to improve it remains a valuable research direction.
@@ 3 Problem Formulation
We formalize NSR as the problem of learning a parameterized symbolic regressor $S_{\boldsymbol{\theta}}$ that maps a numerical dataset $\mathcal{D}$ to a symbolic expression $\hat{e}=S_{\boldsymbol{\theta}}(\mathcal{D})$. The learning algorithm is formulated as minimizing a loss that measures how well $\hat{e}$ matches the ground-truth expression $e^{\ast}$ underlying $\mathcal{D}$. In this section, we specify how synthetic training pairs $(e^{\ast},\mathcal{D})$ are generated in NeSymReS (Biggio et al., 2021), since it is the foundational work underlying our research.
@@ 3.1 Synthetic Expression Distribution
Let $\mathcal{V}=\{x_{1},\dots,x_{d}\}$ be a finite set of variables, $\mathcal{O}_{\mathrm{bin}}$ the binary operators (e.g., ${+,,×,÷}\{$$\,\times$$\div\}$), and $\mathcal{O}_{\mathrm{un}}$ the unary operators (e.g., $\{\sin,\,\cos,\,\log,\,\exp\}$). Denote by $\mathcal{C}=[c_{\min},c_{\max}]\subset\mathbb{R}$ the interval from which numeric constants are drawn. The complete vocabulary for the expression is $\Sigma=\mathcal{V}\cup\mathcal{O}_{\mathrm{bin}}\cup\mathcal{O}_{\mathrm{un}}\cup\mathcal{C}$.
Let $\mathcal{E}$ be an expression space and $p_{\mathcal{E}}$ be the generator of symbolic expressions employed in NeSymReS (Biggio et al., 2021). We also denote by $p_{\mathrm{Tree}}$ the generator of unary-binary trees introduced by Lample and Charton (2019). We write $e^{\ast}\sim p_{\mathcal{E}}$ for the following procedure.
@@ 3.2 Synthetic Dataset Generation
$\mathcal{D}\;=\;\{\,(\mathbf{x}_{i},\,y_{i})\,\}_{i=1}^{n},\quad x_{ij}\;\sim\;\mathcal{U}([x_{\min,j},x_{\max,j}])$
$\text{for}\quad j=1,2,\ldots,d,\quad y_{i}\;=\;e^{\ast}(\mathbf{x}_{i}).$
@@
Note that, in practice, the training dataset is the collection of $e_{\mathrm{templ}}$, and both $e^{\ast}$ and $\mathcal{D}$ are generated dynamically during training. Further details concerning the work of NeSymReS (Biggio et al., 2021) are described in Appendix A.
Figure 1: (Left) Percentage of expressions beyond the training dataset generated by NeSymReS on the not_included dataset. Throughout the training procedure, NeSymReS can hardly generate expressions that are not included in the training data, indicating strong reproduction bias. (Right) NeSymReS exhibits strong fitting performance on the baseline dataset but performs poorly on expressions from the not_included dataset, whose tree structures are absent from the training data. The result indicates the severe effect of reproduction bias on numerical accuracy.
4 Theoretical Analysis on Expression Generation Ability of Transformers
Transformer-based symbolic regression tends to suffer from low performance, particularly when the number of input variables is large. In this section, we explore the theoretical basis of this limitation. Ideally, Transformers should be able to generate tokens in a compositional manner, while maximizing the probability for the final expression to fit the numerical data. However, we show that Transformers inherently lack the capacity to generate expressions compositionally while accounting for their numerical characteristics. We introduce a simplified version of the symbolic regression task and show that Transformers are not expressive enough to solve the task.
We define the last-token prediction problem as the task of predicting the most suitable final token in an otherwise complete mathematical expression. Although predicting the entire optimal expression is NP-hard (Virgolin and Pissis, 2022), this task is much easier since the search space is limited to several leaf tokens. We present a formal definition of this task in the following.
We first introduce $\mathrm{expr}:\Gamma^{\ast}\times\mathbb{R}^{(d+1)*n}\!\longrightarrow\!\mathcal{E}$, a function that maps a token sequence $\mathbf{s}$ to the most appropriate expression $e_{\mathbf{s}}$ that can be represented by $\mathbf{s}$, taking into account the numerical data $\mathcal{D}$. Since the token sequence may contain the placeholder token $C$ representing constants, the mapping is tasked with identifying the optimal values for these constants and transforming the sequence into a corresponding expression tree.
Definition 4.1 (Last-token prediction problem).
Given numerical data $\mathcal{D}$ of $n$ features-value pairs $(\mathbf{x}_{i},y_{i})\in\mathbb{R}^{d}\times\mathbb{R}$, a metric $\mathcal{L}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$, and an incomplete token sequence $\widetilde{\mathbf{s}}$ that forms a prefix representation of an expression when terminated with a leaf token, the last-token prediction problem asks to find a leaf token $u^{\ast}$ such that:
For the analysis, we assume a bounded-depth log-precision Transformer as in (Feng et al., 2023; Merrill and Sabharwal, 2023b, a; Strobl, 2023), a realistic setting where the intermediate computation values of the Transformer are limited to $O(\log k)$ bit precision, with $k$ denoting the maximal length of the input sequence. We now present the theoretical result stating that Transformers with bounded size cannot solve the last-token prediction problem.
Theorem 4.2.
Assume $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$. For any integer $D$ and any polynomial $Q$, there exists a problem size $m$ such that no log-precision Transformer defined in Section G.1 with depth $D$ and hidden dimension $d\leq Q(m)$ can solve $\operatorname{LastTokenPrediction}(m)$.
We show the above theorem by leveraging circuit complexity theory. Specifically, $\mathsf{TC}^{0}$ and $\mathsf{NC}^{1}$ are types of circuit complexity classes, and it is generally conjectured that $\mathsf{TC}^{0}\subsetneq\mathsf{NC}^{1}$. Prior work (Merrill and Sabharwal, 2023b) shows that log-precision Transformers can be simulated with $\mathsf{TC}^{0}$ circuits. We provide a proof for the above theorem by showing that the complexity of the last-token prediction problem is lower bounded by $\mathsf{NC}^{1}$. Detailed specifications of the problem setting and proof of the theorem are provided in Appendix G.
Although the final token of a mathematical expression is arguably the easiest to predict among its components, the above theorem shows that even this seemingly simple task presents substantial difficulties for Transformer models. In general, we find that when Transformers generate tokens via a single forward pass, they lack the capacity to account for numerical consistency, making it questionable whether auto-regressive generation is an appropriate paradigm for symbolic expression generation.
5 Exploring Reproduction Bias in NSR
When generating expressions in an auto-regressive manner, a seemingly appropriate strategy would be to compositionally produce the next token that maximizes accuracy, conditioned on both the previously generated partial expression and the numerical data. However, our theoretical analysis from the last section showed that Transformers lack the ability to do so, meaning that they rely on more coarse-grained strategies. This brings us to the following question: How, in practice, does a transformer generate expressions during inference? In this section, we empirically analyze how expressions are actually generated by NSR models. We demonstrate that NSR models primarily rely on reproduction—that is, they tend to generate expressions by directly copying those seen in the training data.
5.1 Reproduction Bias in Simplified Setting
We first tested how expressions are generated in NeSymReS, which is the method that we mainly focus on in this study. We examined whether the expressions generated by NeSymReS are merely copies from the training data or newly constructed formulas generated compositionally by the model.
Definition of reproduction bias.
We first describe how the novelty of expressions are measured in this section. We introduce $\operatorname{seq}^{-1}:\Gamma^{\ast}\!\longrightarrow\!\mathcal{E}$, which is the inverse function of $\operatorname{seq}$ defined in Section 3.2, and $\operatorname{strip}:\mathcal{E}\!\longrightarrow\!\mathcal{E}$, which is a function that removes all constants from the input expression. We also define $E_{\operatorname{templ}}$ as the set of all $e_{\operatorname{templ}}$ contained in the training dataset.
Definition 5.1 (Reproduction bias).
Given an output token sequence $\mathbf{s}$, let $e=\operatorname{seq}^{-1}(\mathbf{s})$ be the original expression that is represented by $\mathbf{s}$. If $\operatorname{strip}(e)\in E_{\operatorname{templ}}$, we say $\mathbf{s}$ is a reproduction of expressions seen during training.
Remark 5.2.
We have defined and measured reproduction bias based on whether the training dataset contains an expression that is structurally equivalent to the generated one. However, one may argue that we should define and measure reproduction based on functional equivalence; there are many expressions that are structurally different but functionally equivalent (e.g., $x_{1}(x_{1}+x_{2})$ and $x_{1}^{2}+x_{1}x_{2}$), and that such expressions should also be considered as equivalent expressions. To account for both structural and functional equivalence, we choose the operators as described below, so that any pair of structurally different expressions is also functionally different. For a more detailed discussion of the definition of reproduction bias, please refer to Appendix D.
Method.
We constructed a simplified training dataset consisting of 100100K equations. The allowed operator tokens were add, sub, sin, cos, tan, and exp, with up to $5$ independent variables per equation. We then trained a NeSymReS model on this dataset for $1,000$ epochs. The variation of operators was limited due to two reasons. Firstly, the complicated training procedure of NeSymReS, where expressions with operators such as mul or pow are dynamically transformed and presented in different forms across epochs, makes it difficult to judge whether the model’s generated expressions are novel or memorized from training. Therefore, such operators were excluded. Secondly, as explained in Remark 5.2, the choice of operators were restricted in order to account for both structural and functional equivalence when measuring reproduction bias. The dataset size was also kept relatively small due to computational cost and to balance the size of the training data against the size of the search space.